# Neutrino many-body flavor evolution: the full Hamiltonian

### v1

@@ I Introduction
The path towards more realistic many-body studies requires several developments, which include: (i) assessing the impact of using the full many-body Hamiltonian rather than its truncated version; (ii) exploring more general initial pure states (not product states of plane waves), and possibly admixtures of pure states, that more realistically describe the physical system; (iii) studying the dynamics of larger systems and systematically studying the scaling of relevant observables with the number of neutrinos. In this work, we address point (i) above: first, we work out the full neutrino-neutrino Hamiltonian $H_{\nu\nu}$ and set up the framework to implement the time evolution using the occupation number representation. Then we explore the evolution of simple initial states. In this simplified setting, we study the time scales for evolution of flavor and momentum degrees of freedom, and their interplay. While we use a plane wave single-particle basis, we emphasize that in principle any initial state can be built within this formalism, thus allowing one to study point (ii) above.
@@ VI.4 Single neutrino observables
To summarise, the time evolution of $N_{i}^{+}$ and $N_{i}^{-}$ in our model with $N=10$ neutrinos shown in Fig. 9 demonstrate two things. First, non-forward scattering change the flavor evolution of the state by activating momentum modes that are initially unoccupied. It appears that the activation of unoccupied modes accelerate the flavor equilibration. Second, even in such a small system of neutrinos, the initial state that is far from equilibrium can kinetically thermalize. The time scales for flavor and kinatic equilibration appear to be both $t\sim O(1)/\mathcal{E}$, and further studies will be needed to possibly disentangle the two.
@@ VII Conclusions and outlook
Many open questions remain before one can draw definite conclusions about problems of astrophysical interest. We can identify several interesting thrusts for future investigations: (i) Work towards more realistic initial condition within the many-body approach. A key step in this direction is the simulations with larger number of neutrinos ($N$), which will enable a number of interesting investigations. These include studying the scaling of various observables and time scales with $N$; exploring the possible emergence of coherent enhancements, e.g. by considering states with multiple neutrinos within a given solid angle; exploring the effect of spatially non-homogeneous initial conditions; assessing the impact on nucleosynthesis [42] and on the neutrino signal from galactic supernovae. (ii) Implementation on a quantum computer, which requires finding efficient mappings of the full $H_{\nu\nu}$ onto qubit Hamiltonians. (iii) Work towards a comparison with the QKEs, using as common ground the one-body density matrix.

### v2

@@ I Introduction
The path towards more realistic many-body studies requires several developments, which include: (i) assessing the impact of using the full many-body Hamiltonian rather than its truncated version; (ii) exploring more general initial pure states (not product states of plane waves), and possibly admixtures of pure states, that more realistically describe the physical system; (iii) studying the dynamics of larger systems and systematically studying the scaling of relevant observables with the number of neutrinos; including neutrino interactions with other particles. In this work, we address point (i) above: first, we work out the full neutrino-neutrino Hamiltonian $H_{\nu\nu}$ and set up the framework to implement the time evolution using the occupation number representation. Then we explore the evolution of simple initial states. In this simplified setting, we study the time scales for evolution of flavor and momentum degrees of freedom, and their interplay. While we use a plane wave single-particle basis, we emphasize that in principle any initial state can be built within this formalism, thus allowing one to study point (ii) above.
@@ VI.4 Single neutrino observables
To summarise, the time evolution of $N_{i}^{+}$ and $N_{i}^{-}$ in our model with $N=10$ neutrinos shown in Fig. 9 demonstrate two things. First, non-forward scattering change the flavor evolution of the state by activating momentum modes that are initially unoccupied. It appears that the activation of unoccupied modes accelerate the flavor equilibration. Second, even in such a small system of neutrinos, the initial state that is far from equilibrium can kinetically thermalize. The time scales for flavor and kinatic equilibration appear to be both $t\sim O(1)/\mathcal{E}$, and we cannot identify any dependence on the total number of neutrinos $N$. Further studies will be needed to explore whether a separation of the two time scales arise. Kinetic theory suggests that the time scale for momentum evolution should scale as $t_{\rm inc}\sim(G_{F}^{2}T^{5})^{-1}$, in accordance with Fermi’s golden rule, while the time scale for coherent refractive effects should scale as $t_{\rm coh}\sim(G_{F}T^{3})^{-1}$. These estimate rely on the assumption (built into kinetic theory) of interactions localized in space and time, which on the many-body side would require working with wave-packets, which can be built out of our plane waves basis. If one works with plane waves in a box, as shown in Ref. [13] the coherent and incoherent time scales become $t_{\rm coh}\sim 1/({\cal E}N)$ and $t_{\rm inc}\sim 1/({\cal E}\sqrt{N})$, respectively. Therefore, we expect that to disentangle the two will require studying larger systems, with $N\sim O(100)$.
@@ VII Conclusions and outlook
In the many-body approach studied in this paper, several open questions remain before one can draw definite conclusions about problems of astrophysical interest, such as assessing the impact of neutrino flavor evolution on nucleosynthesis [42] and on the neutrino signal from galactic supernovae. We can identify several interesting thrusts for future investigations: (i) A key step is the study of systems with larger number of neutrinos ($N$), which will enable a number of interesting investigations. These include studying the scaling of various observables and time scales with $N$; exploring the possible emergence of coherent enhancements, e.g. by considering initial states with multiple neutrinos within a given solid angle; exploring the effect of spatially non-homogeneous initial conditions (neutrino wavepackets). These studies, besides their intrinsic interest, will also help clarifying the connection between the many-body approach and kinetic theory. (ii) Include neutrino-matter interactions in our formalism. This requires the implementation of additional four-fermion operators in the Hamiltonian, which are however technically simpler than the one studied here because they are linear or bilinear in the neutrino field. (iii) Implementation on a quantum computer, which requires finding efficient mappings of the full $H_{\nu\nu}$ onto qubit Hamiltonians. (iv) Work towards a comparison with the QKEs, using as common ground the one-body density matrix. To quantitatively explore the connection of many-body and QKE approaches will require simulating systems with larger number of neutrinos (see point (i) above and discussion at the end of Section VI.4).